# Scaling Optimal LR Across Token Horizon

### v1

@@ 4.4 Uncertainty Quantification
To ensure that our results are reliable we here consider quantifying the variance of our experiments. We will use two different methodologies. First, following Hoffmann et al. (2022) we will consider the bootstrap method. We consider the 350m model of Table 7, randomly remove 20% of the datapoints and then fit the optimal LR at each token horizon and the scaling law of Equation 3. We repeat this procedure 10 times and then measure the mean and std of the optimal LR and the constants of Equation 3. The results are given in Table 4. We see that the standard deviations are relatively small.
@@
The second methodology we use is simply to rerun a small-scale experiment with multiple seeds. We use the 350m model from Table 7, a 100B token horizon, three learning rates and two seeds. We then estimate the optimal LR via a quadratic fit as in Section 3.1. The results on this experiment is shown in Table 5. We see that there are small differences in the losses, and hence small differences in the estimated optimal LR. The relative std $\sigma/\mu$ is $2.63\!\times\!10^{-2}$, so we can expect the relative error to be a few percent. When using more datapoints random could decrease further due to concentation of measure.
@@
Table 4: We estimate the mean $\mu$ and standard deviation $\sigma$ of the optimal learning rate $LR^{*}$ via bootstrapping. We sample 80% of the data from Figure 2 and then estimate the optimal LR with the procedure from Section 3.1. This bootstrapping precedure is repeated 100 times. We see that the variance is small compared to the mean, the relative error $\sigma/\mu$ is on the order of a few percent. This implies that the uncertainty in our estimates are small.
1 2.940372 2.919948 2.913585 $5.81\!\times\!10^{-4}$
2 2.941199 2.919131 2.912387 $5.76\!\times\!10^{-4}$
3 2.941648 2.920779 2.915190 $5.47\!\times\!10^{-4}$
Table 5: We consider the loss $L$ as a function of the LR. We repeat the experiments of Section 3.1 with a 100B token horizon, a 350m model and multiple seeds. With different seeds we see slightly different final losses, and slightly different estimates of the optimal LR.
@@ 4.5 Effect of Batch Size
It is well known that batch size $B$ affects the optimal LR (Goyal, 2017). While we we primarily focus on the setting of a fixed batch size, we here consider modifying the batch size for the 1.3B model of Table 7. Specifically, we double the batch size to 1m tokens and train for 25B and 50B tokens in total with different LRs. In Figure 10 in the Appendix we show the final validation loss as a function of the LR, and estimate the optimal LR. In Figure 9 we show the optimal LR as a function tokens for the 1.3B model using a batch size of 0.5 or 1m tokens. We see that the optimal LR is higher with larger batch size as expected from Goyal (2017). More importantly we note that the linear fits are roughly parallel. This suggests that the optimal LR depends on token horizon the same way irrespective of the batch size, i.e. that we can factorize Equation 3 as $LR(B,D)=f(B)D^{-\beta}$.
Figure 9: Optimal LR as a function of token horizon for a 1.3B model using a batch size of 0.5m or 1m tokens. Larger batch size implies larger optimal LR as expected. The two lines are roughly parallel, suggesting that the dependence on token horizon is the same irrespective of the batch size.
@@ 5 Related Work
The study of scaling laws for LLMs originated in Kaplan et al. (2020); Henighan et al. (2020) but is still an active area of research. Scaling laws is researched in the context of mainstream topics such post-training Lin et al. (2024); Zhang et al. (2024a); Gao et al. (2023), model architecture Krajewski et al. (2024); Alabdulmohsin et al. (2024); Frantar et al. (2023) and multi-modality Aghajanyan et al. (2023); Cherti et al. (2023). Other areas include inference Sardana & Frankle (2023), multi-linguality Fernandes et al. (2023), synthethic data Dohmatob et al. (2024) and model reuse Wang et al. (2023). There are also more theoretical studies Michaud et al. (2024); Caballero et al. (2022) and applications to recommendation systems Zhang et al. (2024b), game-playing Neumann & Gros (2022) and hyperparameter optimization Kadra et al. (2024).
LR has long been known as an important hyperparameter, and hence there is ample work on how to select it. Goyal (2017) demonstrates that LR needs to be increased when the batch size increases. Larger models are known to require smaller LR, and e.g. Kaplan et al. (2020) suggests the formula $LR(N)\approx 0.003239-0.0001395\log(N)$ for tuning LR as a function of model size. The muP paremetrization (Yang et al., 2022) is a principled method for selecting hyperparameters when the model is scaled, and it will indeed decrease the LR as the width of the model is increased. Validating and extending muP is still an active area of research Lingle (2024); Everett et al. (2024); Noci et al. (2024); Blake et al. (2024). Recently Bi et al. (2024) fits the optimal LR as a function of total compute (which will combine batch size, model size and total training duration) for two small models. Tissue et al. (2024) have provided empirical laws on how the LR schedule affects the validation loss during intermediate steps of training.
@@ 6.1 Limitations
To limit the scope and computational requirements of our study we have intentionally focused on a narrow area – scaling token horizons and changing LR with otherwise fixed LLM recipes. With this limited scope there are naturally many limitations to our study. We have only extended the scaling laws to roughly 1T tokens. Many SOTA LLMs are trained significantly longer, e.g. Llama-3 was trained on 15T tokens (Dubey et al., 2024). Unfortunately, doing ablations experiments at the 15T tokens horizon is not computationally feasible for us. Another limitation is that we haven’t distinguished between repeated and fresh tokens. It is known that one can repeat unique tokens a couple of times with very limited downside (Muennighoff et al., 2024), but exactly how this interacts with optimal LR is left for future studies. Data quality is also known to be very important for LLMs (Abdin et al., 2024), and it likely also interacts with the LR. It is well-known that optimal LR depends on model size, and this study has only scratched the surface of this important topics. Beyond just model size, model architectural modifications like mixture of experts (Shazeer et al., 2017), different attention types (Ainslie et al., 2023), state-space models (Gu & Dao, 2023) could plausibly interact with both the LR and token horizon. There are also multiple hyperparameters which we do not study, which are known to interact with LR. These include batch size (Goyal, 2017) and weight decays (Bjorck et al., 2021), and their full interaction with $LR$ is not clear from our study. At last, while We focus on text-only LLMs there many exciting applications are multimodal LLMs which combine images and textual tokens (Huang et al., 2023). How the LR depends on the number of text tokens and image tokens is an important question left for future work.
6.2 Advice for Practioners
@@
Our experiments show that the optimal LR decreases with token horizon. This necessitates hyperparameter transfer where the optimal LR is found at a shorter horizon and then is modified to be approximately optimal at longer horizons. For practioners with ample compute we recommend finding the best LR at multiple short horizons using the quadratic fitting of Section 3.1. Thereafter the constants in Equation 3 can be found, and the optimal LR at longer horizons can be estimated. For practitioners with less computational resources who are working on larger models (say $>=350m$) we recommend simply using Equation 9 where we have already found $\beta=0.34$ to generalize across architectures. To find the optimal LR $LR^{*}(D_{1})$ at some long horizon $D_{1}$ practitioners can just find the optimal LR $LR^{*}(D_{2})$ at a short horizon $D_{2}$ and then estimate:
$LR^{*}(D_{2})\approx LR^{*}(D_{1})\bigg{(}\frac{D_{2}}{D_{1}}\bigg{)}^{0.34}$ (10)
Since practioners typically find the optimal LR as smaller horizons anyway, this methodoloy has no overhead over current practices.
6.3 Conclusions
@@
We have investigated how learning rate (LR) and token horizon interacts when training LLMs. First, we have show that the optimal LR decreases as the token horizon gets longer. This finding is robust across model sizes. Secondly we have shown that the optimal LR follows reliable scaling laws, and that fitting these allows for hyperparameter transfer accross token horizons. As a case study we have applied our methods to the training of LLama, and have shown evidence which suggests that Llama was training with a LR which was significantly larger than optimal. We argue that this demonstrates how hyperparameter transfer accross token horizons is an imporant and understudied aspect of LLM training. There are many unanswered questions left in this direction, and future studies with additional computational resources can study how token horizons and model size jointly affects optimal LR.
@@ Appendix B Additional Experimental Data
Model size $R^{2}$
2.7b 0.9973
760m 0.9763
350m 0.9607
125m 0.9895
50m 0.9977
Table 8: The $R^{2}$ values of the fits in Figure 4. These are relatively close to 1, indicating a good fit.

### v2

@@ 3.4 Quantifying Variance
To ensure that our results are reliable we here consider quantifying the variance of our experiments. We will use two different methodologies. Firstly we use bootstrapping, following Hoffmann et al. (2022). We consider the 350m model of Table 4, randomly remove 20% of the data points, and then fit the optimal LR at each token horizon and the scaling law of Equation 1. We repeat this procedure 1000 times and then measure the mean and std of the optimal LR and the constants of Equation 1. The results are given in Table 2. We see that the standard deviations are relatively small, which again suggests that the variance in our experimental results is low.
@@
The second methodology we use is simply rerunning a small-scale experiment with multiple seeds. We use the 350m model from Table 4, a 100B token horizon, three learning rates, and two additional seeds. We then estimate the optimal LR via a quadratic fit as in Section 3.1. The results of this experiment are shown in Table 7 in Appendix B. We see that there are small differences in the losses, and hence small differences in the estimated optimal LR. The relative std $\sigma/\mu$ is $2.63\!\times\!10^{-2}$, so we can expect the relative error to be a few percent. When using more data points the variance could further decrease due to concentration of measure.
@@
Table 2: We estimate the mean $\mu$ and standard deviation $\sigma$ of the optimal learning rate $LR^{*}$ via bootstrapping. We sample 80% of the data from Figure 2 and then estimate the optimal LR with the procedure from Section 3.1. This bootstrapping procedure is repeated 1000 times. We see that the variance is small compared to the mean, and the relative error $\sigma/\mu$ is on the order of a few percent. This implies that the uncertainty in our estimates is small.
@@ 3.5 Effect of Batch Size
It is well known that batch size $BS$ affects the optimal LR (Goyal, 2017). While we primarily focus on the setting of a fixed batch size, we here consider modifying the batch size for the 1.3B model of Table 4. Specifically, we double the batch size to 1m tokens and train for 25B and 50B tokens in total with different LRs. In Figure 9 in the Appendix we show the final validation loss as a function of the LR and estimate the optimal LR. In Figure 13 we show the optimal LR as a function of token horizon for the 1.3B model using a batch size of 0.5 or 1m tokens. We see that the optimal LR is higher with a larger batch size as expected. More importantly, we note that the linear fits are roughly parallel. This suggests that the optimal LR depends on the token horizon the same way irrespective of the batch size, i.e. that we can factorize Equation 1 as $LR(BS,D)=f(BS)D^{-\beta}$.
Figure 6: (a) There is a linear relationship between $\log D$ and $\log LR^{*}$ for different model sizes $N$. (b) There is a linear relationship between $\log N$ and $\log LR^{*}$ for different values of dataset size $D$.
4 Scaling Law with respect to model size
So far we have considered Equation 1 with constants fitted independently for each model size. Fully determining the joint scaling properties of the model architecture and token horizon is outside of the scope of this paper – doing so would be computationally infeasible. Nonetheless, we will here provide some preliminary results and discussions regarding a scaling law for both model size $N$ and token horizon $D$. In Figure 6(b) we plot the optimal LR as a function of $N$, and in Figure 6(a) we plot it as a function of $D$. Both curves are straight lines when the axes are logarithmic. This observation motivates the following functional form, which is mathematically derived in Appendix C.
$LR^{*}(N,D)=CN^{-\alpha}D^{-\beta}$ (3)
Here, the constant $C$ is the ideal learning rate for a model of size $1B$ and token horizon $1B$, where the precise numerical value of $C$ may depend on the batch size (see section 3.5), vocabulary size and tokenization. The factor $N^{-\alpha}$ captures the fact that the optimal LR decreases with model size $N$, while the factor $D^{-\beta}$ captures that optimal LR also decreases with token horizon $D$. We now fit Equation 3 to our data for the models of size 760m, 1.3B, and 2.7B, using the 7B model of Section 4.1 as a held-out validation set. We use the huber loss with $\delta=1\times 10^{-3}$ and the BFGS algorithm, similar to Hoffmann et al. (2022). We observe a good fit with RMSE$=2.2\times 10^{-5}$ and a validation $R^{2}$ of $0.978$. The numerical values we find are:
$C\sim 1.55\times 10^{-3},\quad\alpha\sim 0.23,\quad\beta\sim 0.32$ (4)
We depict the results in Figure 7. For smaller models (mainly 50m and 125m) a larger $\beta$ is observed from the experiments. We thus consider Equation 3 to be valid in the regime of large models ($\geq$ 760M parameters).
Figure 7: Fit of Equation 3 compared to the experimental results. The data points for the 7B model of Section 4.1 are excluded at the time of fitting and used as validation data. We have an $R^{2}$ of $0.978$ on this validation data.
4.1 A Case-study on Llama-1
We now consider evaluating if the LRs used by Llama-1 Touvron et al. (2023) are ”correct” according to our scaling laws. To do this we adopt the LLama-1 architecture (RMSnorm, Rope embeddings, and so on) and run small-scale experiments with token horizons 25B, 50B, and 100B and different LRs. Based upon these experiments, we find that values for Equation 1 are $B=8.29\times 10^{-4}$ and $\beta=0.3$ (see Figure 12 in Appendix B). We then extrapolate to find the optimal LR at 1T tokens, which comes out as $1.15\!\times\!10^{-4}$. LLama-1 used $3\!\times\!10^{-4}$, an LR which is too large by a factor of $>2.5$ according to our results. This methodology is visualized in Figure 8. The numerical values used for our predictions are also tabulated in Table 8 in Appendix B, where we also compare to predicting the optimal LR for LLama-1 using our scaling law from Eq. Equation 3. If Llama used the wrong LR, what were the actual effects of this? We provide an estimate on the upper bound for the validation loss penalty between the optimal and used learning rate in Llama-1 using the parabola fit at $D=100B$. $\Delta L=L(D=100B,LR=LR^{*})-L(D=100B,LR=2.6LR^{*})=0.027$. Note, this is an upper bound since the parabolas flatten with longer token horizons
Figure 8: The optimal LR of the LLama-1 7B model as a function of dataset size. We find that $1.15\!\times\!10^{-4}$ is optimal for 1T tokens whereas the Llama-1 paper used $3\!\times\!10^{-4}$.
@@ 5 Related Work
The study of scaling laws for LLMs originated in Kaplan et al. (2020) and it is still an active area of research. Scaling laws are researched in the context post-training (Lin et al., 2024; Zhang et al., 2024a; Gao et al., 2023), model architecture (Krajewski et al., 2024; Alabdulmohsin et al., 2024; Frantar et al., 2023; Wang et al., 2023), multi-modality (Aghajanyan et al., 2023; Cherti et al., 2023), inference (Sardana & Frankle, 2023), data (Dohmatob et al., 2024; Fernandes et al., 2023) and other domains (Zhang et al., 2024b; Neumann & Gros, 2022; Kadra et al., 2024). There are also more theoretical studies (Michaud et al., 2024; Caballero et al., 2022).
LR has long been known as an important hyperparameter, and hence there is ample work on how to select it (Goyal, 2017). Larger models are known to require smaller LR, and e.g. Kaplan et al. (2020) suggests the formula $LR(N)\approx 0.003239-0.0001395\log(N)$ for tuning LR as a function of model size. MuP (Yang et al., 2022) is a principled methodology for selecting LR when the model is scaled, and the method is actively researched (Lingle, 2024; Everett et al., 2024; Noci et al., 2024; Blake et al., 2024). Recently Everett et al. (2024) showed that LR transfer across model size is possible both with different parametrizations and optimizers. However, a limitation in both Yang et al. (2022) and Everett et al. (2024) is the fixed training horizon assumption used in the theoretical derivations. Our work directly explores this limitation. One notable conclusion in Everett et al. (2024) is that extrapolating a LR to larger models may significantly overestimate optimal LR in the compute optimal setting. We reach a different conclusion, the exponent for model size is independent of the token horizon and vice versa. While their results might hold true for other optimizers, a closer look at Table 11 in Everett et al. (2024) shows that for the Adam optimizer, the scaling exponents are similar for the two token horizons they considered. At last, we mention that Bi et al. (2024) recently fits the optimal LR as a function of total compute (which will combine batch size, model size, and total training duration) for two small models.
@@ 6 Discussion
Limitations. To limit the scope and computational requirements of our study we have intentionally focused on a narrow area – scaling token horizons and changing LR with otherwise fixed LLM recipes. With this limited scope, there are naturally many limitations to our study. We have only extended the scaling laws to roughly 800B tokens, while many SOTA LLMs are trained significantly longer (Dubey et al., 2024). It is well-known that optimal LR depends on model size, and this study has only scratched the surface of this important topic. Beyond just model size, model architectural modifications like mixture of experts (Shazeer et al., 2017), different attention types (Ainslie et al., 2023), and state-space models (Gu & Dao, 2023) could plausibly interact with both the LR and token horizon. Other dimensions to consider are additional hyper parameters (Bjorck et al., 2021) and multi-modality (Huang et al., 2023). We leave these topics for future work.
@@
Advice for Practioners. Our experiments show that the optimal LR decreases with the token horizon. This necessitates hyperparameter transfer across token horizons. For practitioners who are working on larger models (say $>=760m$) we recommend simply using Equation 3 where we have already found $\beta=0.32$ to generalize across architectures. To find the optimal LR $LR^{*}(D_{1})$ at some long horizon $D_{1}$ practitioners can just find the optimal LR $LR^{*}(D_{2})$ at a short horizon $D_{2}$ and then estimate:
$LR^{*}(D_{2})\approx LR^{*}(D_{1})\bigg{(}\frac{D_{2}}{D_{1}}\bigg{)}^{-0.32}$ (5)
Since practitioners typically find the optimal LR for smaller horizons anyway during hyperparameter search, this methodology has no overhead over current practices. For practitioners with ample compute resources, we recommend finding the best LR at multiple short horizons using the quadratic fitting of Section 3.1. Thereafter the constants in Equation 1 can be found, and the optimal LR at longer horizons can be estimated. For practitioners using methods that provide zero-shot transfer of optimal LR across width (Yang et al. (2022) and Everett et al. (2024)), we recommend using an adjustment of Equation 3 ($N\propto{n_{layer}}{d^{2}_{model}})$ which yields the formula $LR^{*}(n_{layer},D)=C{n^{-\alpha}_{layer}}D^{-\beta}$. Then $C$ needs to be found using the usual muP sweep for LR with small width, depth and horizon. This allows hyperparameter transfer of LR across depth.
@@
Conclusions. We have investigated how LR and token horizon interact when training LLMs. First, we have shown that the optimal LR decreases as the token horizon gets longer. This finding is robust across model sizes. Secondly, we have shown that the optimal LR follows reliable scaling laws and that fitting these allows for hyperparameter transfer across token horizons. As a case study, we have applied our methods to the training of LLama, and have shown evidence that Llama-1 was trained with a LR which was significantly larger than optimal. We argue thus argue that hyperparameter transfer across token horizons is an understudied aspect of LLM training.
@@ Appendix B Additional Experimental Data
Model size $R^{2}$ $\beta$
2.7b 0.9973 0.4184
1.3b 0.9898 0.3171
760m 0.9763 0.3155
350m 0.9607 0.3799
125m 0.9895 0.6531
50m 0.9977 0.7029
Table 5: The $R^{2}$ values and found $\beta$s of the fits in Figure 4. The $R^{2}$s are relatively close to 1, indicating a good fit. The $\beta$s are relatively similar for models $\geq$ 350m, but are larger for the smallest models.
@@
1 2.940372 2.919948 2.913585 $5.81\!\times\!10^{-4}$
2 2.941199 2.919131 2.912387 $5.76\!\times\!10^{-4}$
3 2.941648 2.920779 2.915190 $5.47\!\times\!10^{-4}$
Table 7: We consider the loss $L$ as a function of the LR. We repeat the experiments of Section 3.1 with a 100B token horizon, a 350m model, and multiple seeds. With different seeds, we see slightly different final losses, and slightly different estimates of the optimal LR.
Figure 13: Optimal LR as a function of token horizon for a 1.3B model using a batch size of 0.5m or 1m tokens. A larger batch size implies that a larger LR is optimal, as expected. The two lines are roughly parallel, suggesting that the dependence on token horizons is the same irrespective of the batch size.
Token Horizon Optimal LR
25B $3.4\times 10^{-4}$
50B $2.8\times 10^{-4}$
100B $2.3\times 10^{-4}$
1T (Using Eq. 1 $1.15\times 10^{-4}$
1T (Using Eq. 3) $1.1\times 10^{-4}$
Table 8: Optimal LR for LLama-1 for different token horizon using the fit in Figure 12. Note that the real Llama-1 used an LR of $3\!\times\!10^{-4}$ which our results suggest is significantly too large.
25B 50B 100B 200B 400B 800B
Optimal LR $1.34\!\times\!10^{-3}$ $1.02\!\times\!10^{-3}$ $6.60\!\times\!10^{-4}$ $4.12\!\times\!10^{-4}$ $2.51\!\times\!10^{-4}$ $1.98\!\times\!10^{-4}$
Predicted LR $4.77\!\times\!10^{-4}$ $3.35\!\times\!10^{-4}$ $2.35\!\times\!10^{-4}$
Ratio 0.864 0.749 0.843
Table 9: Simulated hyperparameter transfer across token horizons for a 125m model. The setup follows Table 9.
$R^{2}$ $\beta$
$\mu$ 0.961 0.384
$\sigma$ $8.85\!\times\!10^{-3}$ $7.78\!\times\!10^{-3}$
$\sigma/\mu$ $9.21\!\times\!10^{-3}$ $2.03\!\times\!10^{-2}$
Table 10: We show the mean $\mu$, standard deviation $\sigma$ and relative deviation $\sigma/|\mu|$ estimated via bootstrapping for two quantities : the $R^{2}$ and $\beta$ in Section 3.4. The relative deviation of $\beta$ is roughly 10%, demonstrating that the uncertainty is reasonably small in our scaling laws.
Appendix C Derivation
$LR^{*}(N,D)=B(N)D^{-\beta(N)}$ (6)
$LR^{*}(N,D)=A(D)N^{-\alpha(D)}$ (7)
The constants $A,\alpha$ may depend on $D$. The key question that arises is how the general notion of model size $N$ can be incorporated into the joint scaling law. Moreover, the scaling law formula from Eq. 7 for constant $D$ has to be representable by Eq. 6. It is anticipated to align with the latter, consisting of distinct power laws, each with specific parameters for different $N$ and $D$ values. Consequently, the objective is to identify a function that fulfills these criteria
$LR^{*}(N,D)=B(N)D^{-\beta(N)}=A(D)N^{-\alpha(D)}$ (8)
$B(N)N^{\alpha}=A(D)D^{\beta}$ (9)
The LHS of Eq. 9 only depends on $N$, whereas the RHS only depends on $D$ so they should both equal some constant, $C$ (this step relies on our proof above that $\alpha,\beta$ are independent of $N,D$), resulting in the functional forms of $A(D),B(N)$
$A(D)=CD^{-\beta},B(N)=CN^{-\alpha}$ (10)
Plugging the functional forms of $B(N)$ we finally get the final functional form for the joint scaling law as in Eq. 3.